Over the past few years the satellite Voyager II has passed by the planets Saturn, Uranus, and Neptune, continually updating information about these planets, including the number of moons for each. Uranus is now believed to have 13 more moons than Neptune. Also, Saturn is now believed to have 2 more than twice the number of moons of Neptune. If the total number of moons for these planets is 47, find the number of moons for each planet.
I'm not sure how to turn that into an equation to solve.
thanks

Answer by checkley71(8403)

U=N+13
S=2N+2
U+S+N=47
(N+13)+(2N+2)+N=47
4N+15=47
4N=47-15
4N=32
N=32/4
N=8 MOONS FOR NEPTUNE.
U=8+13=21 MOONS FOR URANUS.
S=2*8+2=16+2=18 MOONS FOR SATURN.
PROOF:
8+21+18=47
47=47